# Training Small Multimodal Models to Bridge Biomedical Competency Gap: A Case Study in Radiology Imaging

### v2

@@ Abstract
The scaling laws and extraordinary performance of large foundation models motivate the development and utilization of such large models in biomedicine. However, despite early promising results on some biomedical benchmarks, there are still major challenges that need to be addressed before these models can be used in real-world applications. Frontier models such as GPT-4V still have major competency gaps in multimodal capabilities for biomedical applications. Moreover, pragmatic issues such as access, cost, latency, and compliance make it hard for clinicians to use privately-hosted state-of-the-art large models directly on private patient data. In this paper, we explore training open-source small multimodal models (SMMs) to bridge biomedical competency gaps for unmet clinical needs. To maximize data efficiency, we adopt a modular approach by incorporating state-of-the-art pre-trained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med. We conduct a comprehensive study of this approach on radiology imaging. For training, we assemble a large dataset with over 1 million image-text pairs. For evaluation, we propose a clinically driven novel approach using GPT-4 and demonstrate its parity with expert evaluation. We also study grounding qualitatively using attention. For best practice, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B). Training on over 1 million image-text pairs took just two days using a standard 8-A100 cluster, which bodes well for further fine-tuning by clinicians using their own data. LLaVA-Rad is fast and can be run on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications.
@@ Introduction
Figure 1: LLaVA-Rad Overview. A We assemble a large dataset with over 1 million chest X-ray image-text pairs; B The corresponding radiology reports are cleaned and structured using GPT-4; C We adopt a modular approach by incorporating state-of-the-art pre-trained models for individual modalities, and focusing on training lightweight adapters, resulting in LLaVA-Rad which outperforms much larger models like GPT-4V and Med-PaLM M on prior standard metrics (F); For evaluation, we also propose a (E) novel error scoring approach using GPT-4 and demonstrate its parity with expert evaluation and (D) a qualitative visualization of the model’s attention during its generative process.
@@
Foundation models, which are models trained from a massive amount of unlabelled data using self-supervised learning, can be quickly adapted to various downstream tasks using only a small amount of task-specific labeled data [5, 35]. Due to the costly biomedical data annotation [21, 20], foundation models have become a new paradigm in biomedicine by achieving state-of-the-art results on many applications, including medical question answering [46, 35] and medical image classification [4, 3]. Recently, multimodal generative AI has emerged as an exciting frontier in the biomedical domain, expanding the application scope from single-modality to multi-modality (e.g., text and image), such as visual question answering and radiology report generation [28, 46, 43]. While existing models are still largely evaluated on artificial biomedical benchmarks, their promising performance demonstrates their potential in medical decision support.
@@
However, there are still bottlenecks hindering the progress toward using these foundation models in real-world clinical settings. First, sharing patient data with large foundation models hosted on the cloud is subject to privacy and reproducibility concerns [44]. Therefore, clinicians may prefer to run inference and fine-tuning locally. Second, existing state-of-the-art models are often very large and resource-intensive, which make local deployment challenging. Smaller models offer decreased serving costs and latency, which is of particular importance in resource-constrained settings outside of data centers [50], as well as offer a smaller carbon footprint [45]. While such smaller language models [^1] We refer to small language models as a smaller (¡10 billion parameters) yet still performant counterpart to large language models (LMs) have shown success in text-only domains [15, 33, 27], multimodal model performance still lags behind large models [46, 43]. Third, many state-of-the-art models are inaccessible [50], necessitating the development of effective open-source models for biomedicine. Finally, even the best models are still subject to hallucinating false information and there exist limited methods to reliably verify the correctness of model outputs at scale, especially in the specialized field of biomedicine [10].
@@
To bridge this gap between existing medical foundation models and real-world clinical applications, we have developed LLaVA-Rad, a small multimodal model (SMM) that attains state-of-the-art (SOTA) performance in standard radiology imaging tasks. We focus our study on identifying key findings from chest X-ray (CXR) images, the most commonly performed medical imaging examination. Automatically drafting high-quality radiology reports is a challenging, yet clinically meaningful, task which could directly increase radiologist productivity and potentially improve communication and decrease burnout [26]. Existing frontier models such as GPT-4V still have a large competency gap even on such a fundamental medical application. To develop LLaVA-Rad, we adopt a modular approach by incorporating state-of-the-art open-source pretrained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med [28].
@@
For training, we assemble a large dataset comprising 1,034,660 radiology image-report pairs from 8 diverse sources. Some data sources only contain structured labels of key findings, in which case we use GPT-4 to synthesize the report based on the labels.
@@
For evaluation, we report standard metrics such as n-gram-based BLEU and ROUGE, as well as factuality checks based on CheXpert and RadGraph [42, 24]. Additionally, we propose G-Rad, a novel factuality evaluation method based on GPT-4. Compared to existing automatic metrics, we show that G-Rad is more consistent with error quantification by practicing radiologists, thus demonstrating the potential of using GPT-4 for evaluation in a manner that is both scalable and highly relevant to medical practice.
@@
To establish best practice for biomedical multimodal learning, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B) [46]. E.g., LLaVA-Rad outperforms GPT-4V on G-Rad by more than four-fold.
@@
LLaVA-Rad is fast and can be run on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications. LLaVA-Rad training is very efficient, taking just two days on over 1 million image-text pairs using a standard 8-A100 cluster. This means that clinicians can further efficiently fine-tune the model as needed using their private data. By examining the model weights, we found that LLaVA-Rad can ground key regions of abnormalities to generated words in the output report, which signals future opportunities to synergize with latest progress in biomedical segmentation and grounded report generation.
@@
In sum, we develop LLaVA-Rad, a lightweight yet high-performing radiology multimodal model for clinical applications. The promising performance of LLaVA-Rad shows that its underlying modular approach can effectively and efficiently bridge the biomedical competency gap in existing frontier models, enabling clinical access with limited computational resources. GPT-4 is used extensively in data processing, synthetic training data generation, and complex evaluation, demonstrating how frontier LLMs can play a pivotal role in training the next generation of biomedical multimodal models.
@@ Results
LLaVA-Rad represents an emerging paradigm in exploring SMMs, following the proliferation of small language models (SLMs). Despite its compact size, which is over an order of magnitude smaller than prior state-of-the-art models such as Med-PaLM M, LLaVA-Rad attains state-of-the-art performance on standard radiology tasks. This bodes well for potential clinical applications with limited computational resources.

### v3

# Towards a clinically accessible radiology foundation model: open-access and lightweight, with automated evaluation

@@ Abstract
The scaling laws and extraordinary performance of large foundation models motivate the development and utilization of such models in biomedicine. However, despite early promising results on some biomedical benchmarks, there are still major challenges that need to be addressed before these models can be used in real-world clinics. Frontier general-domain models such as GPT-4V still have significant performance gaps in multimodal biomedical applications. More importantly, less-acknowledged pragmatic issues, including accessibility, model cost, and tedious manual evaluation make it hard for clinicians to use state-of-the-art large models directly on private patient data. Here, we explore training open-source small multimodal models (SMMs) to bridge competency gaps for unmet clinical needs in radiology. To maximize data efficiency, we adopt a modular approach by incorporating state-of-the-art pre-trained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med. For training, we assemble a large dataset of over 697 thousand radiology image-text pairs. For evaluation, we propose CheXprompt, a GPT-4-based metric for factuality evaluation, and demonstrate its parity with expert evaluation. For best practice, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B). Moreover, LLaVA-Rad requires only one day to be trained on over 697 thousand image-text pairs using a standard 8-A100 GPU cluster, allowing further fine-tuning by clinicians using their own data. The inference of LLaVA-Rad is fast and can be performed on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications.
@@ Introduction
Foundation models, trained on massive amounts of unlabelled data using self-supervised learning, enable rapid adaptation to various downstream tasks with minimal requirement for task-specific labeled data [5, 36]. Due to the high cost of annotating biomedical data [22, 21], foundation models are poised to become a new paradigm in biomedicine, achieving state-of-the-art results on many applications, including medical question answering [47, 36] and medical image classification [4, 3]. Recently, multimodal generative AI has emerged as an exciting frontier in the biomedical domain, expanding the application scope from single-modality to multi-modality (e.g., text and image), such as visual question answering and radiology report generation [29, 47, 44]. While existing models are still largely evaluated on artificial biomedical benchmarks, their promising performance demonstrates their potential in clinical applications.
@@
However, there are still major bottlenecks hindering the progress toward using foundation models in real-world clinical settings. First, sharing patient data with large foundation models hosted on the cloud is subject to privacy concerns [45]. Therefore, clinicians may prefer to run inference and fine-tuning locally. Second, existing state-of-the-art models are often very large and resource-intensive, which makes local deployment challenging. Smaller models incur smaller carbon footprint [46] and offer reduced serving costs and latency, which is of particular importance in resource-constrained settings outside of data centers [51]. However, while small language models have shown success in text domains [16, 34, 28], small multimodal models still have significant performance gaps compared to larger models [47, 44]. Third, many state-of-the-art models are inaccessible [51], necessitating the development of effective open-source models for biomedicine. Finally, even the best models are still subject to errors such as hallucination and there exist limited methods to reliably verify the correctness of model outputs at scale, especially in the specialized field of biomedicine [10].
@@
To bridge this gap between existing medical foundation models and real-world clinical applications, we have developed LLaVA-Rad, a small multimodal model (SMM) that attains state-of-the-art performance in standard radiology imaging tasks (Fig. 1). We focus our study on identifying key findings from chest X-ray (CXR) images, the most commonly performed medical imaging examination. Automatically drafting high-quality radiology reports is a challenging but clinically meaningful task that could directly increase radiologist productivity and potentially improve communication and decrease burnout [27]. Existing frontier models such as GPT-4V still have a large performance gap even on such a fundamental medical application. To develop LLaVA-Rad, we adopt a modular approach by incorporating state-of-the-art open-source pretrained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med [29].
@@
For training, we assemble a large dataset comprising 697,435 radiology image-report pairs from 7 diverse sources. Some data sources only contain structured labels of key findings, in which case we use GPT-4 to synthesize the report based on the ground-truth labels. We only apply this GPT-4 synthesis to the training data.
@@
For evaluation, we report standard metrics such as $n$-gram-based BLEU and ROUGE, as well as factuality checks based on CheXpert and RadGraph [43, 25]. Additionally, we propose CheXprompt, a factuality evaluation method based on GPT-4. Compared to existing automatic metrics, we show that CheXprompt is more consistent with error quantification by practicing radiologists, thus demonstrating the potential of using GPT-4 for evaluation in a manner that is both scalable and highly relevant to medical practice.
@@
To establish best practices for biomedical multimodal learning, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B) [47].
@@
LLaVA-Rad inference is fast and can be run on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications. In addition, LLaVA-Rad training is also very efficient, taking just one day on over 697 thousand image-text pairs using a standard 8-A100 cluster. This means that clinicians can further efficiently fine-tune the model as needed using their private data. By examining the model weights, we found that LLaVA-Rad can ground key regions of abnormalities to generated words in the output report, which signifiesfuture opportunities to synergize with latest progress in biomedical segmentation and grounded report generation.
@@
In summary, we develop LLaVA-Rad, a lightweight yet high-performing radiology multimodal model for clinical applications. The promising performance of LLaVA-Rad shows that its underlying modular approach can effectively and efficiently bridge the multimodal performance gap in existing frontier models, enabling clinical access with limited computational resources.
@@ Results
LLaVA-Rad represents an emerging paradigm in exploring small multimodal models (SMMs), following the proliferation of small language models (SLMs) (Fig. 1). Despite its compact size, which is over an order of magnitude smaller than prior state-of-the-art models such as Med-PaLM M, LLaVA-Rad attains state-of-the-art performance on standard radiology tasks. This bodes well for potential clinical applications with limited computational resources.